OpenStudy (anonymous):

You are knitting a rectangular scarf. The pattern you have created will result in a scarf that has a length of 60 inches and a width of 4 inches. However, you happen to have enough yarn to cover an area of 480 square inches. You decide to increase the dimensions of the scarf so that all of your yarn will be used. If the increase in the length is 10 times the increase in the width, what will the dimensions of the scarf be?

jimthompson5910 (jim_thompson5910):

Let x = increase in the length and y = increase in the width

jimthompson5910 (jim_thompson5910):

we are told that "the increase in the length is 10 times the increase in the width" so x = 10y

jimthompson5910 (jim_thompson5910):

The original scarf looks like this 

jimthompson5910 (jim_thompson5910):

the new scarf (when the dimensions are increased) looks like this 

jimthompson5910 (jim_thompson5910):

basically you're adding on x and y to the corresponding dimensions (because our definitions allow us to do this)

jimthompson5910 (jim_thompson5910):

Area A = length * width A = LW A = (60+x)(4+y) A = (60+10y)(4+y) .. plug in x = 10y 480 = (60+10y)(4+y) .. plug in A = 480 (this is the total area we want to use up all the yarn)

jimthompson5910 (jim_thompson5910):

so your goal is to solve 480 = (60+10y)(4+y) for y, then use that solution to find x